OpenStudy (xxanntdmxx):

what is the solution to the equation 8-2b = - 2/3(12b + 6) ? b =____ is it -2

OpenStudy (lynfran):

8-2b=-2/3(12b+6) 8-2b=-8b-4 8+4=-8b+2b 12=-6b 12/-6=b -2=b correct